OpenStudy (idealist10):

Find a particular solution of y"-4y'-5y=-6xe^(-x).

OpenStudy (idealist10):

@freckles @CGGURUMANJUNATH

OpenStudy (idealist10):

I just need to know the substitution yp. Any idea?

OpenStudy (freckles):

well probably (At+B)*exp(-x)

OpenStudy (freckles):

that t is suppose to be x

OpenStudy (freckles):

\[y''-4y'-5y=x \text{ if this was the differential equation my guess for } \\ \text {particular solution would be } y_p=Ax+B \\ y''-4y'-5y=e^{-x} \text{ if this was the differential equation my guess for the } \\ \text{ particular solution would be } y_p=Ce^{-x} \\ \text{ so my guess for } y''-4y'-5y=-6xe^{-x} \text{ would be } y_p=(Ax+B)e^{-x}\]

OpenStudy (idealist10):

But why (Ax+B) for x?

OpenStudy (freckles):

because it works... \[y''-4y'-5y=x \\ y_p=Ax+B \\ y_p'=A \\ y_p''=0 \\ 0-4A-5(Ax+B)=x \\ -4A-5Ax-5B=x \\ -4A-5B-5Ax=x \\ \text{ so } A=\frac{-1}{5} \text{ and } -4 \cdot \frac{-1}{5}-5B=0 \implies B=\frac{-1}{5}(-\frac{4}{5})=\frac{4}{25}\]

OpenStudy (freckles):

if you had y''-4y'-5y=x^7 then I would do the particular solution: \[y_p=Ax^7+Bx^6+Cx^5+ Dx^4+Ex^3+Fx^2+Gx+H\]

OpenStudy (idealist10):

Okay, thanks for the hint.

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx this is a good source took at the table and look at the few sentences below that table

OpenStudy (freckles):

"The more complicated functions arise by taking products and sums of the basic kinds of functions"

OpenStudy (freckles):

that is the sentence part I mainly want you to look at

OpenStudy (freckles):

in the table notice we have a polynomial*exponential the polynomial we should a polynomial of the same degree with unknown coefficients*the exponential part and actually they give an example after table exactly like the one you have here (well kind of exactly)